OpenStudy (anonymous):

simplify this: (x^-3y^3)^2/x^5y^6

OpenStudy (anonymous):

(x^(-3)y^(3))^2/(x^(5)y^(6))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To solve this it is as simple as (x^(-3)y^(3))^2 = (x^(-3)y^(3))(x^(-3)y^(3)) = (x^(-3)y^(3)x^(-3)y^(3)) (x^(-3)y^(3)x^(-3)y^(3)) = x^(-6)y^(6) Remember the rule x^(-1) = 1/x so y^(6)/x^(6) / (x^(5)y^(6))/1 = y^(6)/x^(6)y^(6)x^(5) = 1/x^(11)

OpenStudy (anonymous):

y^(6)/x^(6)y^(6)x^(5) = y^(6)y^(-6)/x^(6)x^(5) = y^(0)/x^(6)x^(5) = 1/x^(6)x^(5) remembering a^(0) = 1, where a =/= 0

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

the first part I you just have to remember that x^(-3)*y^(3) is 1 term so when it is squared it is just (x^(-3)y^(3))(x^(-3)y^(3)) as x^(2) = x * x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so I will assume you get it now :)

OpenStudy (anonymous):

what i s the answer?

OpenStudy (anonymous):

I posted it in my first solution

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

ok i see i confused my self lol

OpenStudy (anonymous):

Yeah I wasn't really clear in my math sorry I just wanted to explain the rules as I went along